OpenStudy (abhisar):

Motor cyclist moving with a velocity of 72 km/hr on a flat road takes a turn on the road at a point where radius of curvature of the road is 20 metres. the acceleration due to gravity is \(\sf 10 m/s^2\). In order to avoid skidding, he must not bend with respect to the vertical plane by an angle greater than

OpenStudy (abhisar):

@Abmon98

OpenStudy (abhisar):

yeah i just figured it out by myself ! :)

OpenStudy (abmon98):

\[V=72km/hr \]\[V=72*1000/60*60 \]\[v=20m/s \] m/s \[\tan \theta=v^2/rg\] \[\tan \theta=20*20/20*10\] if you cannot open the link, good to hear that :D

OpenStudy (abhisar):

yes, we will have to consider the optimum condition and take frictional force as zero.

OpenStudy (abhisar):

Thank you for your time though :)

OpenStudy (abmon98):

my pleasure :D

OpenStudy (abhisar):

Here all the centripetal force is due to the sin component of normal force.